A women bought an number of items for $90. She realized that if she had bought 10 more for the same money she would have paid 12 less per item. How many items did she buy?
Answer by Boreal(15235)

x items * 90/x price per item=90
90/(x+10) is $12 less than 90/x
Therefore (90/(x+10))+12=90/x
multiply everything by (x)(x+10)
90x+12x^2+120x)=90x+900, clearing fractions.
12x^2+120x-900=0
divide everything by 12
x^2+10x-75=0
(x+15)(x-5)=0
x=5, only positive number
Paid $18 per item for 5 items
Had she bought 10 more, or 15 altogether, she would have paid $6 per item, which is $12 less.